OpenStudy (anonymous):

find the derivative of f(x) = 3x - 1/x + 2, x = -2

OpenStudy (anonymous):

You can use the quotient rule to figure out what f'(x) is. When you are asked to find the slope at a specific point (x = -2) you just need to evaluate f'(x) at that point.

OpenStudy (anonymous):

@itwildo

OpenStudy (anonymous):

you can also use the sum rule and remember that \[\frac{ 1 }{ x }= x^{-1}\] then the same rule applies as when you differentiate positive powers of x

OpenStudy (anonymous):

Can be rewritten as: \[f(x) =3x-x ^{-1}+2\] Which makes it simple to derivate: \[f'(x) = 3-(-1)x ^{-2} = 3+(1/x^2)\] Solve for x= -2: \[f'(-2)=3+(1/-2^2)=3.25\]

OpenStudy (guptalakshay558):

How Can you find derivative at that point of function at which function is not defined In Above question, as x+2 is in denominator so x=-2 makes it to tend towards infinite so there is not point of slope at x=-2 .. So derivative at X=-2 is also not defined